A yacht sails 18 km on a bearing 324° followed by 6 km on a bearing 132°.
Calculate the distance (in km to the nearest 0.1 km) and bearing (to the nearest degree) of the yacht's final position from its initial position.
distance (km) ≈
(I had 13.3 but i do not think it is correct?)
bearing (°) ≈
THANK YOU

Answer by Alan3354(69443)

A yacht sails 18 km on a bearing 324° followed by 6 km on a bearing 132°.
Calculate the distance (in km to the nearest 0.1 km) and bearing (to the nearest degree) of the yacht's final position from its initial position.
distance (km)
(I had 13.3 but i do not think it is correct?)
It's a triangle with sides of 18 & 6 with and included angle of 12 degs.
---
c^2 = 18^2 + 6^2 - 2*18*6*cos(12)
c^2 =~ 12.195 km
===============================
Angle at the starting point:
sin(C)/c = sin(A)/a
sin(12)/12.195 = sin(A)/6
sin(A) = 0.1022936
A =~ 5.87 degs
bearing = A + 324 = 329.87°